### v3

# Lyapunov exponents and geodesic stability of Schwarzschild black hole in the non-commutative gauge theory of gravity

Abstract
In this paper, we study the stability of geodesic motion for both massive and massless particles using Lyapunov exponents in the non-commutative (NC) Schwarzschild black hole (BH) via the gauge theory of gravity. As a first step, we investigate the both time-like and null radial motion of particles, the mean result in NC geometry shows that the particles take infinity proper time to reach the NC singularity (infinite time affine parameter framework for photons). The proper/coordinate time of Lyapunov exponents and their ratio of time-like geodesic for the circular motion of this black hole shows a new behavior, which describes a new range of stable circular orbits between unstable ones. Then we analyze the circular motion of photons, where the result shows a new photon sphere near the event horizon which is not allowed in the commutative case, and the Lyapunov exponent is expressed in this geometry, where this confirms the instability of the outer photon sphere and the stability of the inner one. Moreover, we studied the effect of noncommutativity on the black hole shadow radius, We found a similarity between the non-commutativity and the mass of a black hole. Then we using experimental data from the event horizon telescope, we show that a noncommutativity parameter of the order of $\Theta^{\text{Phy}}\sim 10^{-32}m$.
Non-commutative geometry, Schwarzschild space-time, Geodesic equation, Black hole shadow, Lyapunove exponents.
I Introduction
The investigation of geodesic motion for test particles around compact celestial bodies is fundamental in astrophysics, offering valuable insights into the nature of neutron stars, black holes (BHs), and other dense astrophysical objects. The trajectory of particles moving around a BH encodes information about the underlying spacetime geometry, which is shaped by the strong gravitational field of the compact object. This study focuses on analyzing how non-commutative (NC) geometry influences the geodesic motion of test particles near a Schwarzschild BH, employing the Lyapunov exponent Lyapunov5 as a diagnostic tool. Two key types of geodesic motion often examined in BH physics are radial and circular trajectories. A wide range of studies has previously explored geodesic motion around BHs, such as Chandr1 ; gibbons1 ; islam1 ; jaklit ; stuchli ; kran1 ; kran2 ; cruz ; hackmann1 ; kagraman ; hackmann2 ; hackmann3 ; levin ; grunau ; belbrun ; barack ; pugliese1 ; pugliese2 ; virb1 ; virb2 ; pradhan ; olivares ; villanueva ; pugl1 ; pugl2 . The classification of stable and unstable orbits sheds light on the spacetime geometry surrounding these objects. The Lyapunov exponent $\lambda$ provides a means to evaluate the stability of equatorial circular geodesics in BH spacetimes. Depending on the nature of the orbits—whether stable, unstable, or marginally stable—the Lyapunov exponent helps in determining their properties. Circular geodesics in a Schwarzschild BH are integrable, meaning they exhibit no chaotic behavior. In particular, unstable orbits have a positive Lyapunov exponent Lyapunov5 , while stable and marginally stable orbits correspond to imaginary and zero values of the exponent, respectively Lyapunov1 ; Lyapunov2 ; Lyapunov3 ; Lyapunov5 ; Lyapunov6 . The importance of geodesic stability has driven numerous studies that utilize Lyapunov exponents to explore various BH spacetimes geostab12 ; geostab11 ; geostab3 ; geostab8 ; geostab9 ; geostab4 ; geostab10 ; geostab6 ; geostab1 ; geostab2 ; geostab7 .
Recent research has investigated how non-commutativity alters geodesic dynamics using multiple approaches, such as modeling the NC matter distribution through Gaussian functions nozari1 ; nozari2 ; kuniyal ; iftikhar ; larranaga ; bhar ; rahaman and Lorentzian profiles NCLD1 ; NCLD2 ; NCLD3 , in addition to the star product formalism as well as Bopp’s transformation ulh1 ; larbi1 and NC gauge gravity abdellah1 ; abdellah4 ; NCGTG2 ; NCGTG3 ; NCGTGmassdeformed1 ; NCGTGmassdeformed2 ; NCGTGmassdeformed3 . In previous work abdellah1 , we identified a new class of particle motion near the event horizon in an NC spacetime for massive test particles. Here, we extend this analysis to massless particles and other motion types, providing a more comprehensive study of geodesic stability. This work contributes to the ongoing exploration of orbital stability using Lyapunov exponents in NC spacetimes geostab14 ; geostab13 ; geostab5 .
The main objective of this study is to explore the stability of geodesic motion in NC spacetime by analyzing the Lyapunov exponent for different types of particles. We aim to understand how quantum gravity effects, introduced by NC geometry, impact orbital behavior. This investigation enhances our knowledge of stability and chaos in deformed spacetimes, with potential implications for observables such as the innermost stable circular orbit and the BH shadow. NC geometry, which is motivated by string theory seiberg1 , is based on the quantization of spacetime, leading to the discretization of gravitational interactions. These quantum gravitational corrections become particularly relevant in the strong-field regions surrounding BHs and should be considered in such analyses.
The coordinate operators in NC spacetime obey the following commutation relation:
$[x^{\mu},x^{\nu}]_{*}=i\Theta^{\mu\nu},$ (1)
where $\Theta^{\mu\nu}$ represents an antisymmetric real matrix that sets the fundamental discretization scale of spacetime. This modification affects the Heisenberg uncertainty principle, leading to a minimum measurable length on the order of the Planck scale:
$\Delta x^{\mu}\Delta x^{\nu}\geq\frac{1}{2}|\Theta^{\mu\nu}|.$ (2)
In the NC framework, the usual function multiplication is replaced by the Moyal star product ”*”, which is defined for two functions $f(x)$ and $g(x)$ as follows:
$(f*g)(x)$ $=f(x)e^{\frac{i}{2}\Theta^{\mu\nu}\overleftarrow{\partial_{\mu}}\overrightarrow{\partial_{\nu}}}g(x)$ (3)
$=f(x)g(x)+\frac{i}{2}\Theta^{\mu\nu}\partial_{\mu}f(x)\partial_{\nu}g(x)-\frac{1}{8}\Theta^{\mu\nu}\Theta^{\rho\eta}\partial_{\mu}\partial_{\rho}f(x)\partial_{\nu}\partial_{\eta}g(x)+\mathcal{O}(\Theta^{3}).$ (4)
In this study, we adopt the NC gauge gravity approach in conjunction with the Seiberg-Witten (SW) map seiberg1 and star products. NC gauge gravity is a candidate extension of general relativity, maintaining action invariance under NC transformations via the SW map action1 ; action2 ; action3 . We utilize this framework to derive the NC Schwarzschild metric, then compute the NC corrections to the effective potential and analyze geodesic motion for different particle types. Our findings suggest that both massive and massless particles are prevented from reaching the NC singularity within a finite proper time or affine parameter, while also revealing the presence of novel stable circular orbits near the event horizon—distinct from those predicted in other NC models. Further, we assess the stability of these orbits through Lyapunov exponent analysis, confirming their stability in NC spacetime. Additionally, we determine NC corrections to the BH shadow radius and use observational data to estimate the NC parameter, yielding a fundamental length close to the Planck scale.
The structure of this paper is as follows: Section II details the NC corrections to the metric field using the star product and SW maps. Section III derives the NC geodesic equations and the effective potential, incorporating second-order corrections in $\Theta$. We examine radial and circular motion for both massive and massless particles, studying orbital stability through the Lyapunov exponent. This is followed by an exploration of NC corrections to the BH shadow radius and an estimation of the NC parameter using observational data. The final section summarizes our key results and conclusions.
II Non-commutative corrections for Schwarzschild black hole
In this section, we briefly present the NC correction to the Schwarzschild metric within the framework of the gauge theory of gravity. Following the methodology outlined in our previous works abdellah1 ; abdellah6 ; abdellah2 , we employ both the Seiberg-Witten (SW) map and the star ($\ast$)-product to construct the NC metric for the Schwarzschild black hole. To obtain the NC gauge potential of gravity, represented by the tetrad fields, we utilize the perturbative expansion of the SW map for the tetrad fields $\hat{e}^{a}_{\mu}(x,\Theta)$. This expansion is carried out as a power series in $\Theta$ up to second order, as derived in Ref. cham1 .
$\hat{e}_{\mu}^{a}=\,$ $e_{\mu}^{a}-\frac{i}{4}\Theta^{\nu\rho}[\omega^{ac}_{\nu}\partial_{\rho}e^{d}_{\mu}+(\partial_{\rho}\omega^{ac}_{\mu}+R^{ac}_{\rho\mu})e^{d}_{\nu}]\eta_{cd}+\frac{1}{32}\Theta^{\nu\rho}\Theta^{\lambda\tau}\left[2\{R_{\tau\nu},R_{\mu\rho}\}^{ab}e^{c}_{\lambda}-\omega^{ab}_{\lambda}(D_{\rho}R_{\tau\mu}^{cd}+\partial_{\rho}R_{\tau\mu}^{cd})e^{m}_{\nu}\eta_{dm}\right.$
$\left.-\{\omega_{\nu},(D_{\rho}R_{\tau\mu}+\partial_{\rho}R_{\tau\mu})\}^{ab}e^{c}_{\lambda}-\partial_{\tau}\{\omega_{\nu},(\partial_{\rho}\omega_{\mu}+R_{\rho\mu})\}^{ab}e^{c}_{\lambda}-\omega^{ab}_{\lambda}\left(\omega^{cd}_{\nu}\partial_{\rho}e^{m}_{\mu}+\left(\partial_{\rho}\omega_{\mu}^{cd}+R_{\rho\mu}^{cd}\right)e^{m}_{\nu}\right)\eta_{dm}\right.$
$\left.+2\partial_{\nu}\omega_{\lambda}^{ab}\partial_{\rho}\partial_{\tau}e^{c}_{\mu}-2\partial_{\rho}\left(\partial_{\tau}\omega_{\mu}^{ab}+R_{\tau\mu}^{ab}\right)\partial_{\nu}e^{c}_{\lambda}-\{\omega_{\nu},(\partial_{\rho}\omega_{\lambda}+R_{\rho\lambda})\}^{ab}\partial_{\tau}e^{c}_{\mu}-\left(\partial_{\tau}\omega_{\mu}^{ab}+R^{ab}_{\tau\mu}\right)\left(\omega^{cd}_{\nu}\partial_{\rho}e^{m}_{\lambda}\right.\right.$
$\left.\left.+\left((\partial_{\rho}\omega^{cd}_{\lambda}+R^{cd}_{\rho\lambda})\right)e^{m}_{\nu}\right)\eta_{dm}\right]\eta_{cb},$ (5)
and the commutative gauge gravity potentials are denoted by $\hat{e}_{a}^{\mu}$ and $\omega^{ab}_{\mu}$, which are the tetrad fields and the spin connection, and:
$\{\alpha,\beta\}^{ab}=\left(\alpha^{ac}\beta^{db}+\beta^{ac}\alpha^{db}\right)\eta_{cd},\quad$ $[\alpha,\beta]^{ab}=\left(\alpha^{ac}\beta^{db}-\beta^{ac}\alpha^{db}\right)\eta_{cd}$ (6)
$D_{\mu}R_{\rho\sigma}^{ab}=\partial_{\mu}R^{ab}_{\rho\sigma}+$ $\left(\omega_{\mu}^{ac}R^{db}_{\rho\sigma}+\omega_{\mu}^{bc}R^{da}_{\rho\sigma}\right)\eta_{cd}.$ (7)
The component $\hat{e}_{a}^{\mu}$ is the of the vierbein $\hat{e}_{\mu}^{a}$ defined as:
$\hat{e}_{\mu}^{b}\hat{e}_{a}^{\mu}=\delta_{a}^{b},\quad\hat{e}_{\mu}^{a}\hat{e}_{a}^{\nu}=\delta_{\mu}^{\nu}\,.$ (8)
In this context, the NC correction to the metric component $\hat{g}_{\mu\nu}$ is obtained according to the following definition chai1 :
$\hat{g}_{\mu\nu}=\frac{1}{2}(\hat{e}_{\mu}^{b}\ast\hat{e}_{\nu b}+\hat{e}_{\nu}^{b}\ast\hat{e}_{\mu b})\,.$ (9)
In this work we choose the following NC matrix $\Theta^{\mu\nu}$:
$\Theta^{\mu\nu}=\left(\begin{matrix}0&0&0&0\\ 0&0&0&\Theta\\ 0&0&0&0\\ 0&-\Theta&0&0\end{matrix}\right),\qquad\mu,\nu=0,1,2,3$ (10)
In order to obtain the NC correction to the metric we follow the same steps in Ref. abdellah1 , and we take the same choose of the general tetrads field:
$\underline{e}_{\mu}^{0}$ $=\left(\begin{array}[]{cccc}\left(1-\frac{2m}{r}\right)^{\frac{1}{2}},&0,&0,&0\end{array}\right),$ (11b)
$\underline{e}_{\mu}^{1}$ $=\left(\begin{array}[]{cccc}0,&\left(1-\frac{2m}{r}\right)^{-\frac{1}{2}}\sin\theta\cos\phi,&r\,\cos\theta\cos\phi,&-r\,\sin\theta\sin\phi\end{array}\right),$ (11d)
$\underline{e}_{\mu}^{2}$ $=\left(\begin{array}[]{cccc}0,&\left(1-\frac{2m}{r}\right)^{-\frac{1}{2}}\sin\theta\sin\phi,&r\,\cos\theta\sin\phi,&r\,\sin\theta\cos\phi\end{array}\right),$ (11f)
$\underline{e}_{\mu}^{3}$ $=\left(\begin{array}[]{cccc}0,&\left(1-\frac{2m}{r}\right)^{-\frac{1}{2}}\cos\theta,&-r\,\sin\theta,&0\end{array}\right).$ (11h)
The NC tetrads $\hat{e}^{a}_{\mu}$ are abdellah1 :
$\hat{e}^{0}_{0}=$ $\left(1-\frac{2m}{r}\right)^{\frac{1}{2}}+\Theta^{2}\Bigg{\{}\frac{m\left(88m^{2}+mr\left(-77+15\sqrt{1-\frac{2m}{r}}\right)-8r^{2}\left(-2+\sqrt{1-\frac{2m}{r}}\right)\right)\sin^{2}\theta}{32r^{5}(1-\frac{2m}{r})^{3/2}}\Bigg{\}}+\mathcal{O}(\Theta^{3})$ (12a)
$\hat{e}^{1}_{1}=$ $\left(1-\frac{2m}{r}\right)^{-\frac{1}{2}}sin\theta cos\phi-\frac{im\,\Theta}{4r^{2}}\left(1-\frac{2m}{r}\right)^{-\frac{3}{2}}sin\theta sin\phi+\Theta^{2}\Bigg{\{}\frac{m\Bigg{(}m\bigg{(}25+2\sqrt{1-\frac{2m}{r}}\bigg{)}r-2\bigg{(}1+\sqrt{1-\frac{2m}{r}}\bigg{)}r^{2}}{64r^{5}\big{(}1+\frac{2m}{r}\big{)}^{5/2}}$
$+\frac{-24m^{2}+\bigg{(}24m^{2}+m\big{(}-29+6\sqrt{1-\frac{2m}{r}}\big{)}\bigg{)}r+2\bigg{(}5-3\sqrt{1-\frac{2m}{r}})r^{2}\bigg{)}\bigg{)}\cos(2\theta)\Bigg{)}}{64r^{5}\big{(}1+\frac{2m}{r}\big{)}^{5/2}}\Bigg{\}}sin\theta cos\phi+\mathcal{O}(\Theta^{3})$ (12b)
$\hat{e}^{2}_{1}=$ $\left(1-\frac{2m}{r}\right)^{-\frac{1}{2}}sin\theta sin\phi+\frac{im\,\Theta}{4r^{2}}\left(1-\frac{2m}{r}\right)^{-\frac{3}{2}}sin\theta cos\phi+\Theta^{2}\Bigg{\{}\frac{m\Bigg{(}m\big{(}25+2\sqrt{1-\frac{2m}{r}}\big{)}r-2\bigg{(}1+\sqrt{1-\frac{2m}{r}}\bigg{)}r^{2}}{64r^{5}\big{(}1+\frac{2m}{r}\big{)}^{5/2}}$
$+\frac{-24m^{2}+\bigg{(}24m^{2}+m\bigg{(}-29+6\sqrt{1-\frac{2m}{r}}\bigg{)}\bigg{)}r+2\bigg{(}5-3\sqrt{1-\frac{2m}{r}})r^{2}\bigg{)}\bigg{)}\cos(2\theta)\Bigg{)}}{64r^{5}\big{(}1+\frac{2m}{r}\big{)}^{5/2}}\Bigg{\}}sin\theta sin\phi+\mathcal{O}(\Theta^{3})$ (12c)
$\hat{e}^{3}_{1}$ $=\left(1-\frac{2m}{r}\right)^{-\frac{1}{2}}cos\theta+\Theta^{2}\Bigg{\{}\frac{m\Bigg{(}-24m^{2}+\bigg{(}29-6\sqrt{1-\frac{2m}{r}}\bigg{)}r-2\bigg{(}5-3\sqrt{1-\frac{2m}{r}}\bigg{)}r^{2}\Bigg{)}}{32r^{5}\big{(}1+\frac{2m}{r}\big{)}^{5/2}}\Bigg{\}}sin^{2}\theta\cos\theta+\mathcal{O}(\Theta^{3})$ (12d)
$\hat{e}^{1}_{2}=$ $r\cos\theta\cos\phi+\frac{i\Theta}{4}\Bigg{(}1-\bigg{(}1-\frac{2m}{r}\bigg{)}^{-\frac{1}{2}}\Bigg{)}\cos\theta\sin\phi+\Theta^{2}\Bigg{\{}\frac{m\Big{(}\Big{(}6-9\sqrt{1-\frac{2m}{r}}\Big{)}r+\Big{(}-2+5\sqrt{1-\frac{2m}{r}}\Big{)}r\cos(2\theta)}{64r^{3}\big{(}1+\frac{2m}{r}\big{)}^{3/2}}$
$+\frac{\bigg{(}18r+\bigg{(}-28+8\sqrt{1-\frac{2m}{r}}\bigg{)}m\bigg{)}\sin^{2}\theta}{64r^{3}\big{(}1+\frac{2m}{r}\big{)}^{3/2}}\Bigg{\}}\cos\theta\cos\phi+\mathcal{O}(\Theta^{3})$ (12e)
$\hat{e}^{2}_{2}=$ $r\cos\theta\sin\phi-\frac{i\Theta}{4}\Bigg{(}1-\bigg{(}1-\frac{2m}{r}\bigg{)}^{-\frac{1}{2}}\Bigg{)}\cos\theta\cos\phi+\Theta^{2}\Bigg{\{}\frac{m\Big{(}\Big{(}6-9\sqrt{1-\frac{2m}{r}}\Big{)}r+\Big{(}-2+5\sqrt{1-\frac{2m}{r}}\Big{)}r\cos(2\theta)}{64r^{3}\big{(}1+\frac{2m}{r}\big{)}^{3/2}}$
$+\frac{\bigg{(}18r+\bigg{(}-28+8\sqrt{1-\frac{2m}{r}}\bigg{)}m\bigg{)}\sin^{2}\theta}{64r^{3}\big{(}1+\frac{2m}{r}\big{)}^{3/2}}\Bigg{\}}cos\theta sin\phi+\mathcal{O}(\Theta^{3})$ (12f)
$\hat{e}^{3}_{2}=$ $-r\sin\theta+\Theta^{2}\Bigg{\{}\frac{m\bigg{(}-2m\bigg{(}1+2\sqrt{1-\frac{2m}{r}}\bigg{)}+\bigg{(}m\bigg{(}-14+4\sqrt{1-\frac{2m}{r}}\bigg{)}+\bigg{(}11-5\sqrt{1-\frac{2m}{r}}\bigg{)}r\bigg{)}\cos(2\theta)}{64r^{3}\big{(}1+\frac{2m}{r}\big{)}^{3/2}}$
$+\frac{r\bigg{(}1+\sqrt{1-\frac{2m}{r}}\bigg{)}\bigg{)}}{64r^{3}\big{(}1+\frac{2m}{r}\big{)}^{3/2}}\Bigg{\}}\sin\theta+\mathcal{O}(\Theta^{3})$ (12g)
$\hat{e}^{1}_{3}=$ $-rsin\theta sin\phi+\frac{i\Theta}{4}\Bigg{(}1-\bigg{(}1-\frac{2m}{r}\bigg{)}^{-\frac{1}{2}}\Bigg{)}\sin\theta\cos\phi+\Theta^{2}\Bigg{\{}\frac{m\bigg{(}-2r\bigg{(}1-\sqrt{1-\frac{2m}{r}}\bigg{)}+\bigg{(}2m\bigg{(}9-17\sqrt{1-\frac{2m}{r}}\bigg{)}\bigg{)}}{32r^{3}\big{(}1+\frac{2m}{r}\big{)}^{3/2}}$
$+\frac{2r-\bigg{(}13-17\sqrt{1-\frac{2m}{r}}\bigg{)}r\bigg{)}\sin^{2}\theta}{32r^{3}\big{(}1+\frac{2m}{r}\big{)}^{3/2}}\Bigg{\}}\sin\theta\sin\phi+\mathcal{O}(\Theta^{3})$ (12h)
$\hat{e}^{2}_{3}=$ $rsin\theta cos\phi+\frac{i\Theta}{4}\Bigg{(}1-\bigg{(}1-\frac{2m}{r}\bigg{)}^{-\frac{1}{2}}\Bigg{)}\sin\theta\sin\phi-\Theta^{2}\Bigg{\{}\frac{m\bigg{(}-2r\bigg{(}1-\sqrt{1-\frac{2m}{r}}\bigg{)}+\bigg{(}2m\bigg{(}9-17\sqrt{1-\frac{2m}{r}}\bigg{)}\bigg{)}}{32r^{3}\big{(}1+\frac{2m}{r}\big{)}^{3/2}}$
$+\frac{2r-\bigg{(}13-17\sqrt{1-\frac{2m}{r}}\bigg{)}r\bigg{)}\sin^{2}\theta}{32r^{3}\big{(}1+\frac{2m}{r}\big{)}^{3/2}}\Bigg{\}}\sin\theta\cos\phi+\mathcal{O}(\Theta^{3}).$ (12i)
Noting that, in the limit of $\Theta=0$, we recover the commutative tetrad fields (11b-11h).
Using the definition in Eq. (9), along with the above NC tetrad fields, and we stop up to the second order in the NC parameter $\Theta$, one can derive the following nonzero components of the deformed metric $\hat{g}_{\mu\nu}$ in the equatorial plane $\theta=\pi/2$:
$-\hat{g}_{tt}=$ $\left(1-\frac{2m}{r}\right)+\left\{\frac{m\left(88m^{2}+mr\left(-77+15\sqrt{1-\frac{2m}{r}}\right)-8r^{2}\left(-2+\sqrt{1-\frac{2m}{r}}\right)\right)}{16r^{4}(-2m+r)}\right\}\Theta^{2}+\mathcal{O}(\Theta^{4})$ (13a)
$\hat{g}_{rr}=$ $\left(1-\frac{2m}{r}\right)^{-1}+\left\{\frac{m\left(12m^{2}+mr\left(-14+\sqrt{1-\frac{2m}{r}}\right)-r^{2}\left(5+\sqrt{1-\frac{2m}{r}}\right)\right)}{8r^{2}(2m-r)^{3}}\right\}\Theta^{2}+\mathcal{O}(\Theta^{4})$ (13b)
$\hat{g}_{\theta\theta}=$ $r^{2}+\left\{\frac{m\left(m\left(10-6\sqrt{1-\frac{2m}{r}}\right)-\frac{8m^{2}}{r}+r\left(-3+5\sqrt{1-\frac{2m}{r}}\right)\right)}{16(-2m+r)^{2}}\right\}\Theta^{2}+\mathcal{O}(\Theta^{4})$ (13c)
$\hat{g}_{\phi\phi}=$ $r^{2}+\left\{\frac{5}{8}-\frac{3}{8}\sqrt{1-\frac{2m}{r}}+\frac{m\left(-17+\frac{5}{\sqrt{1-\frac{2m}{r}}}\right)}{16r}+\frac{m^{2}\sqrt{1-\frac{2m}{r}}}{(-2m+r)^{2}}\right\}\Theta^{2}+\mathcal{O}(\Theta^{4})$ (13d)
It is evident that, in the limit $\Theta\to 0$, the commutative Schwarzschild solution is recovered. A detailed discussion of the properties of the obtained metric can be found in Ref. abdellah1 for the equatorial plane, while the general case is discussed in Ref. abdellah2 .
As observed, this metric exhibits two distinct surfaces, corresponding to the static limit surface and the event horizon, respectively. These surfaces are determined by solving the following equations:
$\hat{g}_{00}=0\,,\quad 1/\hat{g}_{rr}=0.$ (14)
The solution to this equations give us respectively the NC static limit surface and the event horizon of the NC Schwarzschild BH abdellah1 ; abdellah2 :
$r_{sls}^{\mathrm{NC}}=r_{h}\left[1+\left(\frac{\Theta}{r_{h}}\right)\left(\frac{4\sqrt{5}+1}{32\sqrt{5}}\right)+\left(\frac{\Theta}{r_{h}}\right)^{2}\left(\frac{10+\sqrt{5}}{128}\right)\right],$ (15)
and
$r_{h}^{NC}=r_{h}\left[1+\frac{3}{8}\left(\frac{\Theta}{r_{h}}\right)^{2}\right].$ (16)
It is clear that in the limit of $\Theta=0$, we recover the commutative event horizon $r_{h}=2m$.
III Geodetic motion and its stability around a non-commutative Schwarzschild black hole
In this section, we analyze the stability of circular orbits constrained to the equatorial plane (i.e., $\theta=\pi/2$) within the framework of the NC Schwarzschild spacetime. The corresponding spacetime metric is given by:
$d\hat{s}^{2}=\hat{g}_{tt}(r,\Theta)c^{2}dt^{2}+\hat{g}_{rr}(r,\Theta)dr^{2}+\hat{g}_{\phi\phi}(r,\Theta)d\phi^{2}.$ (17)
The motion of a test particle in the NC spacetime is governed by the Lagrangian:
$2\mathcal{L}=\hat{g}_{tt}(r,\Theta)c^{2}\dot{t}^{2}+\hat{g}_{rr}(r,\Theta)\dot{r}^{2}+\hat{g}_{\phi\phi}(r,\Theta)\dot{\phi}^{2},$ (18)
where the dot denotes differentiation with respect to the affine parameter $\tau$.
Applying the Euler-Lagrange equations, the generalized momenta are obtained as:
$p_{t}$ $=c^{2}\hat{g}_{tt}(r,\Theta)\dot{t}=E=\text{constant},$ (19a)
$p_{\phi}$ $=\hat{g}_{\phi\phi}(r,\Theta)\dot{\phi}=l=\text{constant},$ (19b)
$p_{r}$ $=\hat{g}_{rr}(r,\Theta)\dot{r}.$ (19c)
Since the Lagrangian is independent of $t$ and $\phi$, energy ($p_{t}=E$) and angular momentum ($p_{\phi}=l$) are conserved quantities. Consequently, the expressions for $\dot{t}$ and $\dot{\phi}$ become:
$\dot{t}=\frac{E}{c^{2}\hat{g}_{tt}(r,\Theta)},\quad\dot{\phi}=\frac{l}{\hat{g}_{\phi\phi}(r,\Theta)}.$ (20)
Substituting these relations into the spacetime invariant equation $\hat{g}_{\mu\nu}U^{\mu}U^{\nu}\equiv-h$ (where $U^{\mu}=c^{-1}\frac{dx^{\mu}}{d\tau}$ is the four-velocity), we obtain the radial equation:
$\dot{r}^{2}=-\frac{E^{2}}{c^{2}\hat{g}_{tt}(r,\Theta)\hat{g}_{rr}(r,\Theta)}-\frac{1}{\hat{g}_{rr}(r,\Theta)}\left(\frac{l^{2}}{\hat{g}_{\phi\phi}(r,\Theta)}+hc^{2}\right),$ (21)
where $h=1$ for massive particles and $h=0$ for massless particles.
By incorporating the specific metric components, the radial equation simplifies to:
$\dot{r}^{2}+V_{\text{eff}}(r,\Theta)=0,$ (22)
where the effective potential $V_{\text{eff}}(r,\Theta)$ for a test particle in NC spacetime up to second order in the NC parameter $\mathbf{\Theta}$, is given by:
$V_{\text{eff}}(r,\Theta)$ $=\left(1-\frac{2m}{r}\right)\left(\frac{l^{2}}{r^{2}}+hc^{2}\right)-\frac{E^{2}}{c^{2}}+\Theta^{2}\left\{-\frac{l^{2}}{r^{4}}\left(1-\frac{2m}{r}\right)\left(\frac{5}{8}-\frac{3}{8}\sqrt{1-\frac{2m}{r}}+\frac{m\left(-17+\frac{5}{\sqrt{1-\frac{2m}{r}}}\right)}{16r}\right.\right.$
$\left.\left.+\frac{m^{2}\sqrt{1-\frac{2m}{r}}}{(-2m+r)^{2}}\right)+\frac{E^{2}}{c^{2}}\left(\frac{m(64m^{2}+m(-49+13\sqrt{1-\frac{2m}{r}})r+2(13-3\sqrt{1-\frac{2m}{r}})r^{2})}{16r^{5}(1-\frac{2m}{r})^{2}}\right)\right.$
$\left.+\left(\frac{l^{2}}{r^{2}}+hc^{2}\right)\left(\frac{m(12m^{2}+m(-14+\sqrt{1-\frac{2m}{r}})r-(5+\sqrt{1-\frac{2m}{r}}))r^{2}}{8r^{5}(1-\frac{2m}{r})}\right)\right\}+\mathcal{O}(\Theta^{4})$ (23)
In the commutative limit ($\Theta\to 0$), the effective potential reduces to the standard Schwarzschild form:
$V_{\text{eff}}(r,\Theta=0)=\left(1-\frac{2m}{r}\right)\left(\frac{l^{2}}{r^{2}}+hc^{2}\right)-\frac{E^{2}}{c^{2}}.$ (24)
III.1 Radial motion of massive particles
Here, we analyze the radial motion of a massive test particle in the NC Schwarzschild spacetime described by Eq. (17). In this context, we set $\dot{\phi}=0$ and $l=0$, while assuming $h=1$ and $c=1$. Under these conditions, Eq. (22) reduces to:
$\left(\frac{dr}{d\tau}\right)^{2}$ $=E^{2}-\left(1-\frac{2m}{r}\right)-\Theta^{2}\left\{E^{2}\left(\frac{m(64m^{2}+m(-49+13\sqrt{1-\frac{2m}{r}})r+2(13-3\sqrt{1-\frac{2m}{r}})r^{2})}{16r^{5}(1-\frac{2m}{r})^{2}}\right)\right.$
$\quad\left.+\left(\frac{m(12m^{2}+m(-14+\sqrt{1-\frac{2m}{r}})r-(5+\sqrt{1-\frac{2m}{r}}))r^{2}}{8r^{5}(1-\frac{2m}{r})}\right)\right\}+\mathcal{O}(\Theta^{4}).$ (25)
When $\Theta=0$, this equation reduces to the classical radial motion equation of a massive particle in Schwarzschild spacetime.
Next, we consider the free-fall motion of a massive particle in the NC Schwarzschild BH geometry. The particle is initially at rest, i.e., $\dot{r}=0$, at $r=r_{0}$ when $\tau=0$. For the proper time in NC spacetime, we set $E=1$ in Eq. (25) and integrate, yielding:
$\hat{\tau}$ $=-\int_{r_{0}}^{r}\left(\frac{2m}{r^{\prime}}-\Theta^{2}\left\{\left(\frac{m(64m^{2}+m(-49+13\sqrt{1-\frac{2m}{r^{\prime}}})r^{\prime}+2(13-3\sqrt{1-\frac{2m}{r^{\prime}}})r^{\prime 2})}{16r^{\prime 5}(1-\frac{2m}{r^{\prime}})^{2}}\right)\right.\right.$
$\quad\left.\left.+\left(\frac{m(12m^{2}+m(-14+\sqrt{1-\frac{2m}{r^{\prime}}})r^{\prime}-(5+\sqrt{1-\frac{2m}{r^{\prime}}}))r^{\prime 2}}{8r^{\prime 5}(1-\frac{2m}{r^{\prime}})}\right)\right\}\right)^{-1/2}dr^{\prime}.$
$=-\int_{r_{0}}^{r}\sqrt{\frac{2m}{r^{\prime}}}\left(1-\Theta^{2}\frac{r^{\prime}}{2m}\left\{\left(\frac{m(64m^{2}+m(-49+13\sqrt{1-\frac{2m}{r^{\prime}}})r^{\prime}+2(13-3\sqrt{1-\frac{2m}{r^{\prime}}})r^{\prime 2})}{16r^{\prime 5}(1-\frac{2m}{r^{\prime}})^{2}}\right)\right.\right.$
$\quad\left.\left.+\left(\frac{m(12m^{2}+m(-14+\sqrt{1-\frac{2m}{r^{\prime}}})r^{\prime}-(5+\sqrt{1-\frac{2m}{r^{\prime}}}))r^{\prime 2}}{8r^{\prime 5}(1-\frac{2m}{r^{\prime}})}\right)\right\}\right)^{-1/2}dr^{\prime}.$ (26)
we choose:
$\epsilon$ $=\Theta^{2}\frac{r^{\prime}}{2m}\left\{\left(\frac{m(64m^{2}+m(-49+13\sqrt{1-\frac{2m}{r^{\prime}}})r^{\prime}+2(13-3\sqrt{1-\frac{2m}{r^{\prime}}})r^{\prime 2})}{16r^{\prime 5}(1-\frac{2m}{r^{\prime}})^{2}}\right)\right.$
$\quad\left.+\left(\frac{m(12m^{2}+m(-14+\sqrt{1-\frac{2m}{r^{\prime}}})r^{\prime}-(5+\sqrt{1-\frac{2m}{r^{\prime}}}))r^{\prime 2}}{8r^{\prime 5}(1-\frac{2m}{r^{\prime}})}\right)\right\},$ (27)
where $\epsilon$ is a small parameter, due to the smallness of $\Theta$. Here we can use the first-order approximation of $\epsilon$:
$\left(1-\epsilon\right)^{-1/2}\approx 1+\frac{1}{2}\epsilon+\mathcal{O}\left(\epsilon^{2}\right),$ (28)
Using Eq. (28), proper time in equation (26) can be written up to the second order of $\Theta$ as
[…]
After integration, the proper time expression up to the second order in $\Theta$ is:
$\hat{\tau}$ $=\frac{2}{3}\left(\sqrt{\frac{r_{0}^{3}}{2m}}-\sqrt{\frac{r^{3}}{2m}}\right)+\Theta^{2}\left\{\sqrt{\frac{r_{0}}{2m}}\left(\frac{32m^{2}+4m(-52+\sqrt{1-\frac{2m}{r_{0}}})r_{0}+131r_{0}^{2}}{128(2m-r_{0})r_{0}^{2}}\right)\right.$
$\quad\left.-\sqrt{\frac{r}{2m}}\left(\frac{32m^{2}+4m(-52+\sqrt{1-\frac{2m}{r}})r+131r^{2}}{128(2m-r)r^{2}}\right)+\frac{1}{8m}\left(\text{ArcSin}\left(\sqrt{\frac{2m}{r_{0}}}\right)-\text{ArcSin}\left(\sqrt{\frac{2m}{r}}\right)\right)\right.$
$\quad\left.+\frac{67}{256m}\left(\text{ArcTanh}\left(\sqrt{\frac{r_{0}}{2m}}\right)-\text{ArcTanh}\left(\sqrt{\frac{r}{2m}}\right)\right)\right\}.$ (30)
In the limit $\Theta=0$, this expression reduced to the commutative case, recovering the classical result.
Figure 1: Proper time evolution for the free fall of a massive particle in the NC Schwarzschild BH as a function of $r$, with $m=1$ and the initial position $r_{0}=8$.
Fig. 1 illustrates the variation of proper time $\hat{\tau}$ as a function of $r$ during the free fall of a massive test particle into the NC Schwarzschild BH. The NC parameter $\Theta$ significantly influences the proper time near the event horizon, leading to an increase in its duration. The infinite time is required to reach the NC singularity resulting from non-commutativity effects (see Ref. abdellah1 ), and this phenomenon is absent in the commutative case Chandr1 .
Similarly to proper time, the coordinate time $\hat{t}$ for the free fall in the NC Schwarzschild BH can be obtained using (20) and (25) together with $E=1$:
$\hat{t}$ $=-\int_{r_{0}}^{r}\left(\hat{g}_{tt}\right)^{-1}\left(\frac{2m}{r^{\prime}}-\Theta^{2}\left\{\left(\frac{m(64m^{2}+m(-49+13\sqrt{1-\frac{2m}{r^{\prime}}})r^{\prime}+2(13-3\sqrt{1-\frac{2m}{r^{\prime}}})r^{\prime 2})}{16r^{\prime 5}(1-\frac{2m}{r^{\prime}})^{2}}\right)\right.\right.$
$\left.\left.+\left(\frac{m(12m^{2}+m(-14+\sqrt{1-\frac{2m}{r^{\prime}}})r^{\prime}-(5+\sqrt{1-\frac{2m}{r^{\prime}}}))r^{\prime 2}}{8r^{\prime 5}(1-\frac{2m}{r^{\prime}})}\right)\right\}\right)^{-1/2}dr^{\prime}\,,$ (31)
At leading order in $\Theta$, and integration, the coordinate time is given by:
$\hat{t}$ $=\frac{2}{3}\left(\sqrt{\frac{r_{0}^{3}}{2m}}-\sqrt{\frac{r^{3}}{2m}}\right)+4m\left(\sqrt{\frac{r_{0}}{2m}}-\sqrt{\frac{r}{2m}}\right)-4m\left(\text{ArcTanh}\left(\sqrt{\frac{r_{0}}{2m}}\right)-\text{ArcTanh}\left(\sqrt{\frac{r}{2m}}\right)\right)$
$+\Theta^{2}\left\{\sqrt{\frac{r_{0}}{2m}}\left(-\frac{25}{16r_{0}}-\frac{37m}{128(r_{0}-2m)^{2}}+\frac{15\sqrt{1-\frac{2m}{r_{0}}}}{32(r_{0}-2m)^{2}}-\frac{r_{0}\sqrt{1-\frac{2m}{r_{0}}}}{4(r_{0}-2m)^{2}}-\frac{61}{512(r_{0}-2m)}\right)\right.$
$\left.-\sqrt{\frac{r}{2m}}\left(-\frac{25}{16r}-\frac{37m}{128(r-2m)^{2}}+\frac{15\sqrt{1-\frac{2m}{r}}}{32(r-2m)^{2}}-\frac{r\sqrt{1-\frac{2m}{r}}}{4(r-2m)^{2}}-\frac{61}{512(r-2m)}\right)\right.$
$\left.+\frac{1}{4m}\left(\text{ArcSin}\left(\sqrt{\frac{2m}{r_{0}}}\right)-\text{ArcSin}\left(\sqrt{\frac{2m}{r}}\right)\right)+\frac{605}{1024m}\left(\text{ArcTanh}\left(\sqrt{\frac{r_{0}}{2m}}\right)-\text{ArcTanh}\left(\sqrt{\frac{r}{2m}}\right)\right)\right\}.$ (32)
The commutative expression for the coordinate time is recovered by setting $\Theta=0$.
Figure 2: Coordinate time evolution for the free fall of a massive particle in the NC Schwarzschild BH as a function of $r$, with $m=1$ and the initial position $r_{0}=8$.
In Fig. 2, we present the variation of coordinate time $\hat{t}$ during the free fall of a massive test particle in the NC Schwarzschild BH. The effect of non-commutativity become significant near the NC event horizon, where it lead to an increase in the coordinate time required for the particle to approach the NC horizon. This qualitative behavior aligns with the commutative case, as discussed in Chandr1 , but is further modified due to the presence of non-commutativity.
Figure 3: Comparison of proper and coordinate time for the free fall of a massive particle in the NC Schwarzschild BH as a function of $r$, with $m=1$ and the initial position $r_{0}=8$.
Fig. 3 depicts the behavior of both proper time and coordinate time for a massive particle falling toward the NC Schwarzschild BH. These results indicate that non-commutativity prevents a massive particle from reaching the NC singularity within a finite time, in contrast to the commutative case Chandr1 . It is noteworthy that in this approach, the commutative case is recovered in the limit $\Theta=0$, and preserving the expected large-distance commutative behavior while introducing novel effects near the event horizon. This feature is not realized in other approaches rahaman ; bhar ; larranaga . Furthermore, the results highlight the impact of utilizing the gauge theory of gravity in NC spacetime to analyze the geodesic motion of test particles.
III.2 Circular motion of massive particles and Lyapunov exponents
The effective potential governing the dynamics of a massive test particle is given by (23) for $h=1$, $l\neq 0$, and $c=1$, and is analyzed in detail in Ref. abdellah1 .
Figure 4: Effective potential for a massive particle as a function of $r$, for different values of $\Theta$ with fixed parameters: $E=0.998$, $m=1$, and $l=4.2$.
Fig. 4 illustrates the influence of the NC parameter $\Theta$ on the effective potential of a massive test particle. Notably, the introduction of non-commutativity generates a new minimum in the effective potential near the event horizon, regardless of the angular momentum values ($l\geq 0$). This new minimum, located outside the event horizon, corresponds to a stable circular orbit abdellah1 , leading to the existence of multiple stable circular orbits in NC Schwarzschild geometry. In contrast, prior studies on non-commutativity in geodesic motion—whether based on NC matter distributions nozari1 ; nozari2 ; kuniyal ; iftikhar ; larranaga ; bhar ; rahaman ; NCGD1 ; NCGD3 ; NCLD2 ; NCLD3 and on NC gauge theory of gravity via deformed geometry NCGTG2 ; NCGTG3 or via deformed mass NCGTGmassdeformed1 ; NCGTGmassdeformed2 ; NCGTGmassdeformed3 have not predicted these new stable circular orbits near the event horizon.
To assess the stability of these circular orbits, the effective potential must satisfy the following conditions:
$V_{eff}(r,\Theta)=0,\quad\frac{dV_{eff}}{dr}=0.$ (33)
Using these conditions, the leading-order corrections in $\Theta$ for the energy $E_{c}^{2}$ and angular momentum $L_{c}^{2}$ of circular orbits are given by:
$E^{2}_{c}$ $\simeq\frac{(-2m+r_{c})^{2}}{r_{c}(r_{c}-3m)}-\frac{\left(B(r_{c})+F(r_{c})\sqrt{1-\frac{2m}{r_{c}}}\right)\Theta^{2}}{32r_{c}^{4}(-3m+r_{c})^{2}\sqrt{1-\frac{2m}{r_{c}}}}+\mathcal{O}(\Theta^{4}),$ (34)
$L^{2}_{c}$ $\simeq\frac{mr_{c}^{2}}{r_{c}-3m}-\frac{\left(S(r_{c})+Q(r_{c})\sqrt{1-\frac{2m}{r_{c}}}\right)\Theta^{2}}{32r_{c}(-3m+r_{c})^{2}(r_{c}-2m)\sqrt{1-\frac{2m}{r_{c}}}}+\mathcal{O}(\Theta^{4}),$ (35)
where the coefficients are defined as:
$B(r_{c})$ $=-120m^{4}+162m^{3}r_{c}-71m^{2}r_{c}^{2}+4mr_{c}^{3},\quad F(r_{c})=204m^{4}-174m^{3}r_{c}+41m^{2}r_{c}^{2}-4mr_{c}^{3},$ (36a)
$S(r_{c})$ $=-48m^{4}+126m^{3}r_{c}-55m^{2}r_{c}^{2},\quad Q(r_{c})=-120m^{4}+174m^{3}r_{c}-75m^{2}r_{c}^{2}+8mr_{c}^{3}.$ (36b)
In the limit $\Theta\to 0$, these expressions reduce to their commutative counterparts for Schwarzschild spacetime.
Figure 5: Energy $E_{c}^{2}$ (left panel) and angular momentum $L_{c}^{2}$ (right panel) for circular orbits of massive particles in NC Schwarzschild spacetime, as functions of $r_{c}$.
Fig. 5 presents the behavior of $E_{c}^{2}$ and $L_{c}^{2}$ for circular orbits as functions of $r_{c}$, highlighting the impact of the NC parameter $\Theta$. The left panel shows that NC modifications introduce a region near the event horizon where the energy condition becomes unphysical ($E_{c}^{2}<0$). However, beyond this region, $E_{c}^{2}$ transitions into a physical region ($E_{c}^{2}>0$), indicating the emergence of a new innermost stable circular orbit (ISCO) in NC geometry abdellah1 . In the right panel, the influence of $\Theta$ on angular momentum is evident, with positive values corresponding to the newly formed stable circular orbits near the event horizon. When $\Theta=0$, these additional features disappear, restoring the classical Schwarzschild behavior.
III.2.1 Orbital velocity and time period
The orbital velocity of a massive test particle in the NC Schwarzschild spacetime is defined as geostab3 ; geostab5 :
$\hat{\Omega}_{c}=\frac{\dot{\phi}}{\dot{t}}.$ (37)
Utilizing Eqs. (20), (34), and (35), the orbital velocity, accurate to second order in the NC parameter $\Theta$, is given by:
$\hat{\Omega}_{c}=\sqrt{\frac{m}{r_{c}^{3}}+\frac{X(r_{c})+W(r_{c})\sqrt{1-\frac{2m}{r_{c}}}}{32r_{c}^{6}(r_{c}-2m)^{3}}\Theta^{2}},$ (38)
where the functions $X(r_{c})$ and $W(r_{c})$ are defined as:
$X(r_{c})$ $=-1272m^{5}+2304m^{4}r_{c}-1544m^{3}r_{c}^{2}+451m^{2}r_{c}^{3}-48mr_{c}^{4},$
$W(r_{c})$ $=-192m^{4}r_{c}+298m^{3}r_{c}^{2}-137m^{2}r_{c}^{3}+24mr_{c}^{4}.$ (39)
In the limit $\Theta=0$, the commutative Schwarzschild case is recovered geostab3 .
The time period, or orbital times scale of the coordinate time in NC spacetime is given by geostab3 :
$\hat{T}_{\hat{\Omega}}=\frac{2\pi}{\hat{\Omega}_{c}}.$ (40)
Figure 6: The behavior of the NC orbital velocity (left panel) and the NC orbital period (right panel) for a massive test particle.
Fig. 6 illustrates the orbital velocity (left panel) and the orbital time scale (right panel) of a massive test particle in NC Schwarzschild spacetime as functions of $r_{c}$. The effects of non-commutativity become significant near the NC event horizon $r_{h}^{NC}$. Specifically, non-commutativity increases the orbital velocity in the vicinity of $r_{h}^{NC}$, while its impact become negligible at larger orbital radius. Similarly, the orbital period decreases due to non-commutativity, approaching zero near the NC event horizon. This reduction is most pronounced close to $r_{h}^{NC}$ and becomes negligible for $r_{c}\gg 3$, where NC effects vanish.
III.2.2 Lyapunov exponents
To analyze the stability and instability of circular geodesics for a massive test particle in the NC Schwarzschild spacetime, we employ the Lyapunov exponents $\hat{\lambda}$. The stability of an orbit is determined by the nature of $\hat{\lambda}$: an imaginary $\hat{\lambda}$ indicates a stable orbit, whereas a real $\hat{\lambda}$ signifies instability geostab2 ; geostab10 . The proper time ($\hat{\lambda}_{p}$) and coordinate time ($\hat{\lambda}_{c}$) Lyapunov exponents in NC spacetime are given by geostab3 ; geostab2 ; geostab10 ; geostab1 :
$\hat{\lambda}_{p}=\pm\sqrt{\frac{-V_{eff}^{\prime\prime}(r,\Theta)}{2}},\quad\hat{\lambda}_{c}=\pm\sqrt{\frac{-V_{eff}^{\prime\prime}(r,\Theta)}{2\dot{t}^{2}}}.$ (41)
Using the effective potential (23) along with Eqs. (20), (34), and (35), the Lyapunov exponents are expressed as:
$\hat{\lambda}_{p}$ $=\sqrt{-\frac{m(r_{c}-6m)}{(r_{c}-3m)r_{c}^{3}}+\frac{Z(r_{c})+P(r_{c})\sqrt{1-\frac{2m}{r_{c}}}}{32(r_{c}-2m)^{3}r_{c}^{6}(r_{c}-3m)^{2}}\Theta^{2}},$ (42)
$\hat{\lambda}_{c}$ $=\sqrt{-\frac{m(r_{c}-6m)}{r_{c}^{4}}+\frac{Y(r_{c})+N(r_{c})\sqrt{1-\frac{2m}{r_{c}}}}{32(r_{c}-2m)^{3}r_{c}^{7}}\Theta^{2}}.$ (43)
The functions $Z(r_{c})$, $P(r_{c})$, $Y(r_{c})$, and $N(r_{c})$ are given by:
$Z(r_{c})$ $=16416m^{7}-38544m^{6}r_{c}+34992m^{5}r_{c}^{2}-16116m^{4}r_{c}^{3}+4076m^{3}r_{c}^{4}-551m^{2}r_{c}^{5}+32mr_{c}^{6},$ (44a)
$P(r_{c})$ $=3564m^{6}r_{c}-8058m^{5}r_{c}^{2}+7070m^{4}r_{c}^{3}-2910m^{3}r_{c}^{4}+517m^{2}r_{c}^{5}-24mr_{c}^{6},$ (44b)
$Y(r_{c})$ $=8880m^{6}-16568m^{5}r_{c}+11200m^{4}r_{c}^{2}-3470m^{3}r_{c}^{3}+515m^{2}r_{c}^{4}-32mr_{c}^{5},$ (44c)
$N(r_{c})$ $=1668m^{5}r_{c}-2870m^{4}r_{c}^{2}+1808m^{3}r_{c}^{3}-473m^{2}r_{c}^{4}+24mr_{c}^{5}.$ (44d)
In the limit $\Theta=0$, the commutative Schwarzschild case is recovered geostab1 ; geostab10 .
Figure 7: Proper time (left panel) and coordinate time (right panel) Lyapunov exponents for a massive particle as functions of the circular orbit radius $r_{c}$.
Fig. 7 illustrates the proper time (left panel) and coordinate time (right panel) Lyapunov exponents for a massive test particle in NC spacetime. The coordinate time Lyapunov exponent $\hat{\lambda}_{c}$ is real in the region $r_{h}^{NC}<r_{c}\leq r_{outer}^{ISCO}$, indicating unstable orbits near the event horizon. As the orbit radius $r_{c}$ increases, the instability decreases, and reaching zero at the outer ISCO, $r_{outer}^{ISCO}$. The proper time Lyapunov exponent $\hat{\lambda}_{p}$ exhibits a distinct behavior: instability (real values) is observed in two regions, $r_{h}^{NC}<r_{c}\leq r_{inner}^{ISCO}$ and $r_{c}^{uns}<r_{c}\leq r_{outer}^{ISCO}$. Between these intervals, in the range $r_{inner}^{ISCO}\leq r_{c}\leq r_{c}^{uns}$, $\hat{\lambda}_{p}$ becomes imaginary, indicating stability of the circular orbit. These results confirm the existence of new stable circular orbits near the event horizon, consistent with previous findings abdellah1 .
The ratio of the proper time to coordinate time Lyapunov exponents is given by:
$\frac{\hat{\lambda}_{p}}{\hat{\lambda}_{c}}=\sqrt{\frac{r_{c}}{r_{c}-3m}+\frac{Y(r_{c})(3m-r_{c})-Z(r_{c})+(N(r_{c})(3m-r_{c})-P(r_{c}))\sqrt{1-\frac{2m}{r_{c}}}}{32m(r_{c}-2m)^{3}(r_{c}-3m)^{2}(r_{c}-6m)r_{c}^{2}}\Theta^{2}}.$ (45)
In the limit $\Theta=0$, the commutative case is recovered geostab1 ; geostab10 .
Figure 8: The ratio of proper time to coordinate time Lyapunov exponents for a massive test particle in NC Schwarzschild spacetime as a function of $r_{c}$.
Fig. 8 illustrates the ratio $\hat{\lambda}_{p}/\hat{\lambda}_{c}$ as a function of $r_{c}$ in NC spacetime. The ratio remains real in two distinct regions, separated by an interval where it becomes imaginary, reflecting the behavior observed in the proper time Lyapunov exponent. The analysis of proper time and coordinate time Lyapunov exponents, along with their ratio, reveals a novel feature in the dynamics of circular motion around an NC Schwarzschild BH. Specifically, it predicted a new range of stable circular orbits located between unstable ones.
III.3 Radial motion of massless particles
The radial motion of massless particles in the NC Schwarzschild spacetime can be derived using the effective potential (23) by setting $h=0$, $c=0$, and $l=0$. This leads to the equation governing radial motion:
$\left(\frac{dr}{d\tau}\right)^{2}$ $=E^{2}-\Theta^{2}\left\{E^{2}\frac{m\big{(}64m^{2}+m(-49+13\sqrt{1-\frac{2m}{r}})r+2(13-3\sqrt{1-\frac{2m}{r}})r^{2}\big{)}}{16r^{5}\left(1-\frac{2m}{r}\right)^{2}}\right\}+\mathcal{O}(\Theta^{4}),$ (46)
where the commutative case is recovered when $\Theta=0$.
For a photon freely falling toward the NC Schwarzschild BH and emitted from a radial position $r=r_{0}$ with energy $E=1$, the affine parameter $\tau$ and coordinate time $\hat{t}$ are given by:
[…]
By evaluating these integrals to leading order in $\Theta$, the results are:
$\hat{\tau}$ $=r_{0}-r+\Theta^{2}\bigg{\{}25\ln\left(\frac{r_{0}(r-2m)}{r(r_{0}-2m)}\right)-\frac{1}{128m}\left[\frac{32m^{2}+15mr_{0}+8r_{0}^{2}}{r_{0}^{2}}-\frac{32m^{2}+15mr+8r^{2}}{r^{2}}\right]$
$\quad+\frac{1}{128m}\bigg{[}\left(1-\frac{2m}{r}\right)^{-\frac{1}{2}}-\left(1-\frac{2m}{r_{0}}\right)^{-\frac{1}{2}}-14\big{(}\sqrt{1-\frac{2m}{r}}-\sqrt{1-\frac{2m}{r_{0}}}\big{)}$
$\quad-\frac{13}{3}\big{(}(1-\frac{2m}{r})^{\frac{3}{2}}-(1-\frac{2m}{r_{0}})^{\frac{3}{2}}\big{)}-\frac{35}{2}\frac{r_{0}(r-2m)-r(r_{0}-2m)}{(r_{0}-2m)(r-2m)}\bigg{]}\bigg{\}},$ (49)
$\hat{t}$ $=r_{0}-r+2m\ln\left(\frac{r_{0}-2m}{r-2m}\right)+\Theta^{2}\bigg{\{}-\frac{224m^{2}-201mr_{0}+63r_{0}^{2}}{128r_{0}(r_{0}-2m)^{2}}+\frac{224m^{2}-201mr+63r^{2}}{128r(r-2m)^{2}}$
$\quad-\frac{(17m^{2}-24mr_{0}+8r_{0}^{2})\sqrt{1-\frac{2m}{r_{0}}}}{m(r_{0}-2m)^{2}}+\frac{(17m^{2}-24mr+8r^{2})\sqrt{1-\frac{2m}{r}}}{m(r-2m)^{2}}-\frac{63}{m}\ln\left(\frac{r_{0}(r-2m)}{r(r_{0}-2m)}\right)\bigg{\}}.$ (50)
The commutative limits of the affine parameter and coordinate time are recovered by setting $\Theta=0$ Chandr1 .
Figure 9: Variation of the affine parameter $\hat{\tau}$ and coordinate time $\hat{t}$ along a radial null geodesic of a photon in NC Schwarzschild spacetime as a function of $r$.
Fig. 9 illustrates the behavior of $\hat{\tau}$ and $\hat{t}$ for a photon initially at $r=r_{0}$ falling toward the NC Schwarzschild BH. Notably, in the affine parameter framework of NC spacetime, photons require an infinite affine parameter to reach the NC singularity, in contrast to the commutative case Chandr1 ; radial1 . The same observation applies to coordinate time, where the photon cannot reach the singularity within a finite time. This behavior is consistent with the motion of massive particles and highlights the role of non-commutativity in preventing singularity encounters within a finite temporal or affine framework. Additionally, for massless particles, our analysis predicts a novel behavior in free-fall motion within NC spacetime using the NC gauge theory of gravity. This differs from the approach in Refs. rahaman ; bhar , which alters the commutative behavior and modifies the underlying physics of the spacetime geometry. In contrast, non-commutativity should act as a quantum correction to the metric, reducing to the commutative case when this correction vanishes. While using an NC matter distribution can recover the commutative free-fall motion, as demonstrated in Ref. kuniyal , it does not introduce any new predictions for free-fall dynamics. However, our approach based on the NC gauge theory of gravity preserves the commutative solution while predicting new macroscopic effects of non-commutativity.
III.4 Circular motion of massless particles and Lyapunov exponents
This section examines the null geodesic motion of photons around an NC Schwarzschild black hole (BH). The effective potential for a photon in this spacetime is obtained from (23) by setting $h=0$, $c=1$ and $l\neq 0$.
Figure 10: Behavior of the effective potential for a massless particle with varying non-commutative parameter $\Theta$ and fixed values of $E=0.998$, $m=1$, $l=4.2$ (left panel), and with varying angular momentum $l$ and fixed $E=0.998$, $m=1$, $\Theta=0.2$ (right panel).
Fig. 10 illustrates the influence of the NC parameter $\Theta$ (left panel) and the angular momentum $l$ (right panel) on the effective potential of a massless test particle (photon). As shown in the left panel, introducing non-commutativity leads to the emergence of a new minimum in the effective potential close to the event horizon. This phenomenon is analogous to what is observed for massive test particles (see SubSec. III.2). However, for massless particles, only two extrema of the effective potential exist outside the event horizon, allowing the interpretation of the new minimum as a stable circular orbit. This results in the formation of a stable photon sphere in this spacetime. Additionally, the position of this new minimum shifts farther from the event horizon as $\Theta$ increases. Moreover, the influence of non-commutativity becomes negligible in regions significantly distant from the event horizon ($r\gg r_{h}^{NC}$). The right panel of Fig. 10 reveals that in NC spacetime, the effective potential exhibits extremal points only for angular momentum values that satisfy a specific condition. This contrasts with the commutative case, where the effective potential always possesses a maximum, irrespective of the angular momentum. For $\Theta=0.2$, the effective potential exhibits two extrema when the angular momentum exceeds a critical value, $l^{\text{crit}}\geq 2.55$. These critical values depend on the NC parameter $\Theta$. In conclusion, NC spacetime consistently supports the existence of two photon spheres: an inner stable photon sphere and an outer unstable photon sphere. This behavior differs significantly from the commutative case and highlights the unique role of NC geometry in modifying the geodesic structure of spacetime. Furthermore, these results are novel and distinct from previous works kuniyal ; bhar ; rahaman ; NCLD2 ; NCGTG2 ; NCGTG3 ; NCGTGmassdeformed1 ; NCGTGmassdeformed2 ; NCGTGmassdeformed3 , which could not predict the existence of a new stable photon sphere near the event horizon.
III.4.1 Null circular orbit and photon sphere
The conditions for the circular motion of photons in NC Schwarzschild spacetime are analogous to those for massive particles, as given in (33). By utilizing the effective potential (23) with $h=0$, the relationship between the photon’s energy $E_{c}$ and angular momentum $L_{c}$ can be expressed as:
$\frac{1}{D_{c}}=\frac{E_{c}}{L_{c}}=\sqrt{\frac{r_{c}-2m}{r_{c}^{3}}+\left(\frac{J(r_{c})+G(r_{c})\sqrt{1-\frac{2m}{r_{c}}}}{16r_{c}^{6}(r_{c}-2m)}\right)\Theta^{2}},$ (51)
where $D_{c}$ is the impact parameter, and the functions $J(r_{c})$ and $G(r_{c})$ are defined as:
$J(r_{c})=156m^{3}-185m^{2}r_{c}+73mr_{c}^{2}-10r_{c}^{3},\quad G(r_{c})=33m^{2}r_{c}-37mr_{c}^{2}+6r_{c}^{3}.$ (52)
Table 1: Numerical values for the unstable $r_{\text{uns}}$ and the new stable circular orbit $r_{\text{sta}}$ of the photon sphere in NC spacetime for different $\Theta$ values, with $E=0.998$, $l=4.2$, and $m=1$.
$\Theta$ 0 0.10 0.15 0.20 0.25 0.30
$r_{\text{sta}}$ (new) 2.18400 2.24624 2.30383 2.35865 2.41183
$r_{\text{uns}}$ 3.00000 2.99782 2.99504 2.99101 2.98559 2.97855
Table 1 presents numerical solutions of (33), illustrating the dependence of the stable and unstable photon sphere radii on $\Theta$. For $3.29\leq l\leq 4.48$ (where this condition depends on $0\leq\Theta\leq 0.3$), the stable circular orbit radius increases with $\Theta$, while the unstable orbit radius decreases. However, for $l>4.48$, both orbit radii increase as $\Theta$ grows.
Figure 11: Behavior of the photon sphere radius in NC spacetime. Stable and unstable circular orbits are shown as functions of $\Theta$ for fixed $l=4.2$, $E=1$, $m=1$ (two left panels), and as functions of $l$ for fixed $\Theta=0.3$, $E=1$, $m=1$ (two right panels).
The left panels of Fig. 11 demonstrate that as $\Theta$ increases, the radius of the stable circular orbit grows, while the radius of the unstable orbit decreases. In the NC framework, the unstable photon orbit has a smaller radius than in the commutative case for $l<4.48$, whereas for $l\geq 4.48$, the unstable orbit radius exceeds the commutative value. This divergence arises from modifications to the gravitational field induced by non-commutativity. The right panels of Fig. 11 illustrate the dependence of circular orbit radii on angular momentum $l$. For a fixed $\Theta=0.3$, the stable circular orbit radius decreases as $l$ increases, while the unstable orbit radius grows. This behavior reflects the distinctive interplay between angular momentum and the NC parameter in shaping the geodesic structure of spacetime.
NC SBH Singularity NC event horizonStable photon sphereUnstable photon sphere
Figure 12: The schematic picture of the stable/unstable photon sphere around the NC SBH.
III.4.2 Lyapunov exponents
In the following, we analyze the stability and instability of the two photon spheres in NC spacetime using the Lyapunov exponent. As is well known, for null circular orbits of photons, proper time does not exist, leaving only the coordinate time Lyapunov exponent for analysis.
We begin by calculating the angular frequency and the period in coordinate time, as given by equations (37) and (40), respectively:
$\hat{\Omega}_{c}^{\text{Null}}=\frac{1}{D_{c}},\quad\hat{T}_{\hat{\Omega}}^{\text{Null}}=2\pi D_{c}.$ (53)
Next, the coordinate time Lyapunov exponent for photon orbits around the NC Schwarzschild BH is derived as:
$\hat{\lambda}_{c}^{\text{Null}}=\sqrt{-\frac{3((-4m+r_{c})(-2m+r_{c}))}{r_{c}^{4}}+\left(\frac{O(r_{c})+I(r_{c})\sqrt{1-\frac{2m}{r_{c}}}}{32(r_{c}-2m)^{2}r_{c}^{7}}\right)\Theta^{2}},$ (54)
where
$O(r_{c})$ $=-19488m^{5}+39000m^{4}r_{c}-29076m^{3}r_{c}^{2}+9958m^{2}r_{c}^{3}-1716mr_{c}^{4}+140r_{c}^{5},$ (55a)
$I(r_{c})$ $=-3171m^{4}r_{c}+5773m^{3}r_{c}^{2}-3908m^{2}r_{c}^{3}+1080mr_{c}^{4}-84r_{c}^{5}.$ (55b)
Figure 13: Coordinate time Lyapunov exponent for massless particles as a function of $r_{c}$ (left panel) and the NC parameter $\Theta$ (right panel).
Fig. 13 illustrates the behavior of the coordinate time Lyapunov exponent for photons as a function of $r_{c}$ (left panel) and $\Theta$ (right panel). The left panel shows that instability initially increases to a peak (decreasing with higher $\Theta$), then decreases as $r_{c}$ increases, eventually vanishing for $r_{c}\geq 4$. In NC spacetime, orbits near the event horizon exhibit reduced instability compared to the commutative case, whereas at larger distances from the event horizon, this trend reverses. Moreover, while instability begins at the event horizon in commutative spacetime, NC geometry shifts it outward, allowing for the presence of stable orbits near the event horizon. This observation is consistent with the effective potential analysis for massless test particles, confirming the existence of a stable photon sphere around the NC Schwarzschild BH. Similar behavior has been reported in other NC geometries, as discussed in Ref. geostab5 , where orbits within the event horizon cannot form a stable photon sphere. Examining the dependence of circular orbit radii $r_{c}$ on $\Theta$ (right panel) reveals that the orbit at $r_{c}=3$ remains unstable in NC spacetime, with increasing $\Theta$ amplifying this instability. In contrast, the new circular orbits emerging near the event horizon in NC spacetime exhibit reduced instability, which decreases as $\Theta$ increases and eventually stabilizes. This suggests the existence of stable circular orbits near the event horizon in NC spacetime.
The ratio of the Lyapunov exponent to the angular frequency in coordinate time ($\frac{\hat{\lambda}_{c}^{\text{Null}}}{\hat{\Omega}_{c}^{\text{Null}}}$) defines the instability exponent of null circular orbits in NC spacetime:
$\frac{\hat{\lambda}_{c}^{\text{Null}}}{\hat{\Omega}_{c}^{\text{Null}}}=\sqrt{\frac{3(4m-r_{c})}{r_{c}}+\left(\frac{O(r_{c})+6J(r_{c})(8m^{2}-6mr_{c}+r_{c}^{2})+(I(r_{c})+6G(r_{c})(8m^{2}-6mr_{c}+r_{c}^{2}))\sqrt{1-\frac{2m}{r_{c}}}}{32(r_{c}-2m)^{3}r_{c}^{4}}\right)\Theta^{2}}.$ (56)
Figure 14: Ratio of the Lyapunov exponent to the angular frequency in coordinate time as a function of $r_{c}$.
Fig. 14 illustrates the variation of $\hat{\lambda}_{c}^{\text{Null}}/\hat{\Omega}_{c}^{\text{Null}}$ with $r_{c}$, showing behavior similar to that in Fig. 13. Unlike the commutative case, NC spacetime allows for stable orbits near the event horizon, in contrast to the results in Ref. geostab5 , which confirm the stability of the new photon sphere.
Finally, the critical exponent $\hat{\gamma}$ for null circular orbits is defined as the ratio of the Lyapunov timescale $\hat{T}_{\hat{\lambda}}=\frac{2\pi}{\hat{\lambda}}$ to the orbital timescale $\hat{T}_{\hat{\Omega}}=\frac{2\pi}{\hat{\Omega}}$ geostab11 ; geostab3 ; geostab12 :
$\hat{\gamma}^{\text{Null}}=\frac{\hat{T}_{\hat{\lambda}}^{\text{Null}}}{\hat{T}_{\hat{\Omega}}^{\text{Null}}}=\frac{1}{2\pi\sqrt{\frac{3(4m-r_{c})}{r_{c}}+\left(\frac{O(r_{c})+6J(r_{c})(8m^{2}-6mr_{c}+r_{c}^{2})+(I(r_{c})+6G(r_{c})(8m^{2}-6mr_{c}+r_{c}^{2}))\sqrt{1-\frac{2m}{r_{c}}}}{32(r_{c}-2m)^{3}r_{c}^{4}}\right)\Theta^{2}}}.$ (57)
In the limit $\Theta=0$, we recover the commutative case.
Table 2: Critical exponent $\hat{\gamma}^{\text{Null}}$ for unstable orbits $r_{\text{uns}}$ and stable orbits $r_{\text{sta}}$ of the photon sphere in NC spacetime for different values of $\Theta$, with $m=1$.
$\Theta$ $r_{c}=3$ $r_{c}=2.4$ $r_{c}=2.3$ $r_{c}=2.2$
$\hat{\gamma}^{\text{Null}}$ 0.0 0.159155
0.1 0.158439 0.117291 0.121293 0.251639
0.2 0.156349 0.136225 0.328429 $0-0.0662942i$
0.3 0.153042 0.210585 $0-0.10626i$ $0-0.0397438i$
0.4 0.148746 $0-0.216564i$ $0-0.0665877i$ $0-0.0288573i$
The numerical values of the critical exponent $\hat{\gamma}^{\text{Null}}$ for photons in various circular orbits $r_{c}$ and different values of the non-commutative (NC) parameter $\Theta$ are summarized in Table 2. This table highlights the presence of both real and imaginary values of $\hat{\gamma}^{\text{Null}}$. Real positive values indicate unstable circular orbits geostab11 , whereas imaginary values correspond to stable circular orbits. For the photon sphere orbit in the Schwarzschild BH spacetime at $r_{c}=3$, instability persists even in the NC framework. According to the definition in Eq. (57), this implies that the Lyapunov timescale $\hat{T}_{\hat{\lambda}}$ is shorter than the orbital timescale $\hat{T}_{\hat{\Omega}}$, i.e., $\hat{T}_{\hat{\lambda}}<\hat{T}_{\hat{\Omega}}$ geostab12 . A similar behavior is observed for other values of $r_{c}$ in Table 2, where real values of $\hat{\gamma}^{\text{Null}}$ consistently signify instability. Furthermore, the impact of the NC parameter $\Theta$ on $\hat{\gamma}^{\text{Null}}$ is evident. As $\Theta$ increases, the values of $\hat{\gamma}^{\text{Null}}$ decrease for $r_{c}=3$, indicating increasing instability. The smallest values of $\hat{\gamma}^{\text{Null}}$ correspond to stronger Lyapunov instability geostab3 (see the right panel of Fig. 13). On the other hand, the imaginary values of $\hat{\gamma}^{\text{Null}}$, which denote stable circular orbits, align with the new minima in the effective potential observed for $l=5.2$ (values comparable to those in Table 1). This demonstrates the emergence of a new stable photon sphere in NC spacetime, contrasting with the instability of the external photon sphere in both commutative and NC scenarios.
III.4.3 Black hole shadow
In the equatorial plane ($\theta=\pi/2$), the radius of the BH shadow, $\hat{R}_{\text{shadow}}$, is directly related to the impact parameter $D_{c}$ by the relation $\hat{R}_{\text{shadow}}=D_{c}\big{|}_{r=r_{\text{ps}}}$, where $r^{\text{uns}}_{\text{ps}}$ represents the radius of the unstable photon sphere. This radius can be determined by solving the equation $\frac{dV_{\text{eff}}}{dr}=0$. To leading order in the NC parameter $\Theta$ and the BH mass $m$, the solution is given by:
$r^{\text{uns}}_{\text{ps}}=3m-\left(\frac{-38+540D_{c}^{2}m^{2}}{288m}\right)\Theta^{2}\,.$ (58)
It is important to note that the radius of the stable photon sphere in NC spacetime can only be obtained numerically (see Table 2). Using the above expression, the shadow radius of the NC Schwarzschild BH can be expressed as:
$\hat{R}_{\text{shadow}}=3\sqrt{3}m\left(1+\frac{(1+2\sqrt{3})}{72m^{2}}\Theta^{2}\right)\,,$ (59)
where the corresponding commutative case is recovered by setting $\Theta=0$.
Figure 15: The shadow of the Schwarzschild BH in NC spacetime for different values of the NC parameter $\Theta$ (left panel) with $m=1$, and for varying BH mass $m$ (right panel) with $\Theta=0.4$.
Fig. 15 illustrates the shadow of the NC Schwarzschild BH for different values of the NC parameter $\Theta$ (left panel) and the BH mass $m$ (right panel). As seen in the left panel, the shadow radius increases with increasing $\Theta$. Similarly, the right panel shows that for a fixed NC parameter, the shadow radius grows as the BH mass $m$ increases. This behavior suggests that the effects of non-commutativity and BH mass on the shadow radius are analogous. The NC parameter influences the spacetime geometry in a way that mimics an increase in mass, effectively strengthening the gravitational field. This conclusion is consistent with our earlier results abdellah1 . Furthermore, we observe that the shadow radius of the NC Schwarzschild BH is larger and exhibits a slight increase with increasing $\Theta$, in contrast to the findings in Hassanabadi2 , where it was reported to increase significantly with increasing $\Theta$.
Armed with the above result and considering the constraint on the NC parameter—namely, that the NC correction must be smaller than the precision of the experimental measurement we choose as an implementation of the NC shadow radius of the Schwarzschild BH. Specifically, we consider a typical primordial BH at the final stage of inflation with a mass of $GM\sim 5\times 10^{-5}m$ karimabadi1 and a volume of $r\sim 1.5\times 10^{-4}m$. Moreover, the precision of the experimental measurement, according to ehc , is approximately $\sim 0.1$. Consequently, the upper limit of the NC parameter can be determined as:
$\Bigg{|}\frac{R_{\text{shadow}}-\hat{R}_{\text{shadow}}}{R_{\text{shadow}}}\Bigg{|}=\sqrt{3}\frac{(1+2\sqrt{3})}{24m}\Theta^{2}\leq 0.1\,.$ (60)
According to Ref. PLANCK1 , the physical distance measured at any time after inflation can be obtained by multiplying our results by the square of the cosmological scale factor, $a=10^{-29}$. This is due to our choice of the space-space NC matrix. By applying the above constraint to our system, we find that the NC parameter satisfies:
$\Theta^{\text{Phy}}=\sqrt{a^{2}\Theta^{2}}\leq 3.94\times 10^{-32}\,\text{m}.$ (61)
It is evident that the obtained result is close to the Planck scale and is consistent with the bounds derived in our previous works, where similar constraints were obtained through experimental tests of gravity in the context of the NC gauge theory of gravity abdellah1 ; abdellah4 or by employing the Bopp’s shift transformation larbi1 ; mirza ; rome1 . These studies suggest a bound in the range of $(10^{-34}-10^{-31})$ m. Remarkably, in this work, the estimation of the NC parameter is directly related to the accuracy of the experimental measurement of the BH shadow radius. This implies that as the precision of the measurement improves, the estimation of the NC parameter can become more refined. A similar observation was made in Ref. abdellah4 , where variations in experimental accuracy led to different values of the NC parameter.
IV Conclusions
In this study, we examined the geodesic motion of both massive and massless test particles in the NC Schwarzschild spacetime. Our approach was based on modifying the Schwarzschild BH metric using the NC gauge theory of gravity abdellah1 . We computed corrections to the effective potential up to second order in $\Theta$ and obtained modifications to both the affine parameter and the coordinate time in the NC Schwarzschild BH. Our results indicate that due to the influence of non-commutativity, both massive and massless particles require an infinite amount of time to reach either the NC singularity or the NC event horizon.
Furthermore, we analyzed the stability of geodesic motion by utilizing the Lyapunov exponent to assess the stability of circular orbits for both massive and massless particles. For massive particles, we determined the angular frequency $\hat{\Omega}_{c}$, the time period $\hat{T}_{c}$, and both the proper-time and coordinate-time Lyapunov exponents, $\hat{\lambda}_{p}$ and $\hat{\lambda}_{c}$, respectively. Notably, the proper-time Lyapunov exponent revealed the presence of a stable orbital region between two unstable regions near the event horizon, which aligns with previous findings abdellah1 . For massless particles, our analysis of the effective potential uncovered a novel type of motion near the event horizon, indicating the presence of a stable circular orbit. As a result, a second photon sphere emerges in NC spacetime. Consequently, the NC Schwarzschild BH features two distinct photon spheres: a stable inner photon sphere and an unstable outer photon sphere. By employing the coordinate-time Lyapunov exponent $\hat{\lambda}_{c}^{\text{Null}}$, we confirmed the instability of the outer photon sphere, while the inner photon sphere remains stable—a feature unique to NC spacetime. Additionally, we computed the critical exponent $\hat{\gamma}^{\text{Null}}$ to further evaluate the instability of null circular orbits in the NC Schwarzschild BH.
Lastly, we explored the impact of non-commutativity on the BH shadow. Our results suggest a resemblance between the effects of mass and non-commutativity, implying that non-commutativity strengthens the gravitational field of the BH. Key insights from our analysis include the connection between newly identified stable circular orbits near the event horizon and non-commutative effects, as well as an estimated upper bound on the NC parameter, which is approximately $\Theta^{\text{Phy}}\sim 10^{-32}$ m.
Acknowledgements.
This work is supported by project B00L02UN050120230003, Univ. Batna 1, Algeria.